# Mitigating “Epistemic Debt” in Generative AI-Scaffolded Novice Programming using Metacognitive Scripts

### v1

@@ Abstract.
The democratization of Large Language Models (LLMs) has given rise to “Vibe Coding”, a workflow where novice programmers prioritize semantic intent over syntactic implementation. While this lowers barriers to entry, we hypothesize that without pedagogical guardrails, it is fundamentally misaligned with cognitive skill acquisition. Drawing on Kirschner’s (2026) distinction between Cognitive Offloading and Cognitive Outsourcing, we argue that unrestricted AI encourages novices to outsource the Intrinsic Cognitive Load required for schema formation, rather than merely offloading Extraneous Load. This accumulation of “Epistemic Debt” creates “Fragile Experts” — those whose high functional utility is masked by critically low corrective competence.
To quantify and mitigate this debt, we conducted a between-subjects experiment ($N=78$) using a custom Cursor IDE plugin backed by Claude 3.5 Sonnet. Participants were recruited via Prolific and UserInterviews.com to represent “AI-Native” learners. We compared three conditions: Manual (Control), Unrestricted AI (Outsourcing), and Scaffolded AI (Offloading). The Scaffolded condition utilized a novel “Explanation Gate,” leveraging a real-time LLM-as-a-Judge framework to enforce a “Teach-Back” protocol before generated code could be integrated.
Results reveal a “Collapse of Competence”: while Unrestricted AI users matched the productivity of the Scaffolded group ($p<.001$ vs. Manual), they suffered a 77% failure rate in a subsequent AI-Blackout maintenance task, compared to only 39% in the Scaffolded group. Qualitative analysis suggests that successful vibe coders naturally engage in self-scaffolding, treating the AI as a consultant rather than a contractor. We discuss the implications for the long-term maintainability of AI-generated software and propose that future learning systems must enforce Metacognitive Friction to prevent the mass production of unmaintainable code.
@@
copyright: rightsretainedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 13th ACM Conference on Learning at Scale; June 2026; Location TBDisbn: 978-1-4503-XXXX-X/26/06ccs: Social and professional topics Computing educationccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
A prevailing narrative in the technology sector suggests that software engineering as a discipline is nearing its end. Industry leaders and commentators have argued that with the advent of advanced Generative AI (GenAI), “English is the new programming language” (World Governments Summit, 2024; Karpathy, 2023), implying that the barrier to creating software has effectively dissolved. The assumption is that if AI can generate functional code from high-level natural language prompts—a practice colloquially known as “Vibe Coding” (Karpathy, 2024)—then the necessity for human expertise in syntax and logic is obsolete. Indeed, the long march of software engineering has moved toward abstracting away low-level details—from assembly programming to providing expressive power through language definitions closer to human natural language, such as Python. In that trajectory, treating human natural language as the next programming abstraction seems inevitable. However, from the standpoint of learning and the long-term maintainability of software, we posit a contrasting view.
While GenAI has undoubtedly democratized the creation of software, we hypothesize that unrestricted reliance on these tools, particularly in educational contexts, is engineering a latent crisis. If the next generation of developers learns to orchestrate “vibes” (intent) without understanding the underlying mechanics, we are not witnessing the democratization of engineering, but the accumulation of Epistemic Debt.
@@
Epistemic Debt describes the accumulation of functional software artifacts that the user “owns” legally but does not “own” cognitively. Unlike technical debt, which resides in the codebase, epistemic debt resides in the mind of the developer. When the inevitable abstraction leak occurs (Spolsky, 2002)—whether through API deprecation, hallucinated vulnerabilities, or scaling failures—the “Vibe Coder” lacks the mental model required to intervene. If this debt is not serviced during the learning process, the field faces a risk of a “Collapse of Competence,” where the workforce is replete with “Fragile Experts” who are skilled at adopting tools (high utility) but incapable of maintaining the systems they build (low comprehension).
@@
Early hypotheses and evidence observing the over-reliance on GenAI in educational contexts have already started to question whether GenAI acts as a tool or is becoming a crutch (Balzan et al., 2026). According to a CDT national poll, 70% of teachers worry AI might weaken critical thinking and research skills, and more than half of students say they feel less connected to teachers when using AI tools (Laird et al., 2025). Surveys regarding ChatGPT use just a year after its unveiling already raised fears that “learners may rely too heavily on the model” (Kasneci et al., 2023). Furthermore, “diminished epistemic vigilance, superficial learning and emotional dependence on AI” (Yan et al., 2025), as well as metacognitive laziness (Fan et al., 2025), have been observed in specific studies. Indeed, students themselves report cognitive offloading and over-reliance on GenAI tools (Wei et al., 2025). A useful metaphor to understand this phenomenon was posited by Kirschner (2026): the distinction between cognitive offloading and cognitive outsourcing, founded in decades of cognitive and learning science research. On that basis, it is already possible to posit that our hypothesis, pointing to the accumulation of epistemic debt from the over-reliance on GenAI tools and the resulting competence collapse, will hold. Consequently, we must go a step further and ask whether this breakdown is inevitable.
@@
Can we design AI interactions and interventions that allow novices to leverage the speed and productivity benefits of GenAI tools without sacrificing the depth of their learning? In other words, can we arrive at a best-of-both-worlds scenario where high functional utility is achieved while learners retain enough expertise to produce comparable corrective competence?
@@
To that end, we once again turn to learning science research, drawing inspiration from Teachable Agents (Biswas et al., 2005) to design an “Explanation Gate” intervention. This intervention requires novice learners, in a scaffolded setting, to teach-back what they have learned and understood to their AI chatbot before being allowed to use the generated code. We hypothesize that this introduction of “Metacognitive Friction” can restore learner comprehension without significantly degrading productivity. Our research questions are as follows:
@@
  • RQ1 (Functional Utility & Velocity): To what extent does the introduction of automated metacognitive friction (the “Explanation Gate”) impact the immediate functional utility and development velocity of novice programmers compared to unrestricted GenAI usage?
@@
  • RQ3 (Interactional Stances): What qualitative differences in human-AI interaction strategies (e.g., “Consultant” vs. “Contractor” stances) correlate with the retention of corrective competence in unrestricted GenAI environments?
@@ 2. Theoretical Framework
Our investigation is grounded in a synthesis of three complementary theories of learning: Cognitive Load Theory (specifically the distinction between Offloading and Outsourcing) (Sweller, 1988), the Script Theory of Guidance (SToG) (Fischer et al., 2013), and the “Desirable Difficulties” framework (Bjork, 1994). We argue that the interaction between novice programmers and Large Language Models (LLMs) represents a novel pedagogical context where traditional scaffolding metaphors break down, requiring a re-evaluation of how automated assistance supports, or subverts, schema acquisition.
@@
A critical error in current EdTech discourse is the conflation of efficiency (productivity) with learning (schema construction). To dismantle this, we rely on Sweller’s Cognitive Load Theory (Sweller, 1988), specifically the interplay between three types of load:
@@
    Germane Load: The mental effort dedicated to processing information into Long-Term Memory (LTM) schemas.
Expert developers use tools to minimize Extraneous Load, freeing up working memory for high-level problem solving (Germane Load). Kirschner (Kirschner, 2026) defines this as Cognitive Offloading. However, novices lack the schemas to distinguish between syntax (extraneous) and logic (intrinsic). When a novice uses GenAI to Vibe Code, the tool handles both the syntax and the underlying logic. This transition from Offloading to Cognitive Outsourcing allows the learner to bypass the Intrinsic Load entirely (Kirschner, 2026). Without engaging with Intrinsic Load, Germane processing cannot occur, and the learner accumulates epistemic debt, a state where functional output exists without the corresponding mental models to maintain it.
@@ 2.2. Overscripting
Fischer’s Script Theory of Guidance (SToG) posits that external scripts (prompts, roles, instructions) are necessary to help learners coordinate complex tasks (Fischer et al., 2013). Ideally, these scripts provide a temporary structure that is gradually “faded” as the learner’s internal scripts (schemas) mature (Wecker and Fischer, 2010).
We argue that the “Copy-Paste” AI workflow represents a pathological case of Overscripting (Dillenbourg, 2002). Because standard LLMs provide complete, high-fidelity solutions instantly, they act as a rigid script that replaces, rather than regulates, the learner’s cognitive activity. Unlike human tutors who withhold answers to prompt reflection, an unrestricted LLM resolves the problem space immediately, denying the learner the “productive failure” necessary for transfer (Kapur, 2008).
To counteract this, we introduce the concept of Metacognitive Friction. This design paradigm deliberately slows down the interaction loop, forcing the learner to engage in generative processing. This aligns with Bjork’s concept of Desirable Difficulties (Bjork, 1994), specifically the “Generation Effect,” which posits that long-term retention is improved when learners must actively retrieve and reformulate knowledge rather than passively consuming it.
@@ 2.3. Scalable Assessment as the Gatekeeper
Historically, the Teachable Agent paradigm, where students learn by teaching a computer agent, has been shown to be highly effective for deep learning (Biswas et al., 2005). However, verifying that a student has truly “taught” the agent (rather than gaming the system (Baker, 2005)) traditionally required human Teaching Assistants (TAs), limiting the scalability of such interventions.
Recent advancements in LLM-Based Assessment allow us to automate this verification at scale. By leveraging an “LLM-as-a-Judge” framework (Choma et al., 2025), we can convert the Teachable Agent model into a real-time “Explanation Gate.” This system acts as a semantic validity check, ensuring that the learner’s vibe (intent) is backed by sufficient syntactic understanding before code integration is permitted. This effectively closes the loop on Outsourcing, forcing the learner to convert the AI’s output back into their own internal representation before moving forward.
@@ 3.2. The Coding Environment
  • Plugin Architecture: We developed a custom VS Code extension, VibeCheck, which acted as a middleman between the LLM Sidebar and the editor. As illustrated in Figure 1, the extension implemented two core logic layers:
@@
      Rule Enforcement Engine: A background listener that intercepted input events from two sources: user keystrokes/clipboard and the LLM Sidebar’s “Apply” signal. It routed actions based on the experimental condition—blocking external pastes for the Manual group, or intercepting the merge signal for the Scaffolded group.
@@
      Explanation Gate Manager (The Friction Loop): A state machine responsible for the intervention. Upon receiving a halt signal from the Enforcement Engine, this manager rendered a blocking modal overlay. It orchestrated an asynchronous loop with the Judge API, preventing the code merge until the Judge returned a “Pass” evaluation, effectively closing the feedback loop shown in the diagram.
@@
      Telemetry Logger: Recorded every keystroke, chat interaction, and state change with millisecond-precision timestamps for post-hoc analysis.

### v2

@@ Abstract.
The democratization of Large Language Models (LLMs) has given rise to vibe coding, a workflow where novice programmers prioritize semantic intent over syntactic implementation. While this lowers barriers to entry, we hypothesize that without pedagogical guardrails, it is fundamentally misaligned with cognitive skill acquisition. Drawing on Kirschner’s (Kirschner, 2026) distinction between cognitive offloading and cognitive outsourcing, we argue that unrestricted AI encourages novices to outsource the intrinsic cognitive load required for schema formation, rather than merely offloading extraneous load. This accumulation of epistemic debt creates fragile experts: developers whose high functional utility masks critically low corrective competence.
To quantify and mitigate this debt, we conducted a between-subjects experiment ($N=78$) using a custom Cursor IDE plugin backed by Claude 3.5 Sonnet. Participants were recruited via Prolific and UserInterviews.com to represent AI-native learners. We compared three conditions: manual (control), unrestricted AI (outsourcing), and scaffolded AI (offloading). The scaffolded condition utilized a novel Explanation Gate, leveraging a real-time LLM-as-a-Judge framework to enforce a teach-back protocol before generated code could be integrated.
Results reveal a collapse of competence: both AI groups significantly outperformed the manual control on functional utility ($p<.001$) and did not differ from each other ($p=.64$), yet unrestricted AI users suffered a 77% failure rate in a subsequent 30-minute AI-blackout maintenance task, compared to only 39% in the scaffolded group. Qualitative analysis suggests that successful vibe coders naturally engage in self-scaffolding, treating the AI as a consultant rather than a contractor. We discuss the implications for the long-term maintainability of AI-generated software and propose that future learning systems must enforce metacognitive friction to prevent the mass production of unmaintainable code.  [^1] Replication package (extension source, task suite, grading harness, and study protocol): https://github.com/sreecharansankaranarayanan/vibecheck
@@
copyright: rightsretainedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 13th ACM Conference on Learning at Scale; June 2026; Seoul, South Koreaisbn: 978-1-4503-XXXX-X/26/06ccs: Social and professional topics Computing educationccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
A prevailing narrative in the technology sector suggests that software engineering as a discipline is nearing its end. Industry leaders and commentators have argued that with the advent of advanced Generative AI (GenAI), “English is the new programming language” (World Governments Summit, 2024; Karpathy, 2023), implying that the barrier to creating software has effectively dissolved. The assumption is that if AI can generate functional code from high-level natural language prompts (a practice colloquially known as vibe coding) (Karpathy, 2024), then the necessity for human expertise in syntax and logic is obsolete. Indeed, the long march of software engineering has been toward abstracting away low-level details, starting from assembly language toward expressive, human-readable languages such as Python. In that trajectory, treating human natural language as the next programming abstraction seems inevitable. However, from the standpoint of learning and the long-term maintainability of software, we posit a contrasting view.
The rapid adoption of LLM-assisted programming has surfaced a critical vulnerability that extends far beyond the classroom. Industry practitioners are increasingly warning that the velocity of AI code generation is masking a severe decline in developer comprehension, a phenomenon Hall (Hall, 2026) describes as the “hidden cost of AI speed.” This growing crisis was a central focus at the ICSE 2026 panel on “Technical Debt in the AI Era” (Avgeriou and Ozkaya, 2026), signaling a paradigm shift in how maintainability is measured. While traditional software engineering has long managed technical debt, the AI era necessitates a focus on epistemic debt, the widening gap between the complexity of a system and the human developer’s cognitive grasp of it (Ionescu et al., 2020). While GenAI has undoubtedly democratized the creation of software, we hypothesize that unrestricted reliance on these tools is engineering a latent crisis. Yan et al. (Yan et al., 2026) find that routine AI usage progressively weakens the intellectual habits required for rigorous problem-solving, accelerating the accumulation of epistemic debt.
@@
Epistemic debt describes the accumulation of functional software artifacts that the user “owns” legally but does not “own” cognitively. Unlike technical debt, which resides in the codebase, epistemic debt resides in the mind of the developer. When the inevitable abstraction leak occurs (Spolsky, 2002), whether through API deprecation, hallucinated vulnerabilities, or scaling failures, the vibe coder lacks the mental model required to intervene. Dell’Acqua et al. (Dell’Acqua et al., 2026) provide corporate-scale evidence of this dynamic: in a field experiment with 758 BCG consultants, AI assistance boosted quality by 32% and speed by 25% on tasks within AI’s capability frontier, yet produced a 19-percentage-point drop in correctness on tasks outside that frontier, because workers had “fallen asleep at the wheel,” over-relying on AI output in domains where it could not be trusted. Critically, the boundary between inside and outside the frontier is opaque to the worker; they cannot tell, in advance, which tasks will expose the limits of AI competence. This is precisely the condition that makes epistemic debt dangerous at scale. If this debt is not serviced during the learning process, the field faces a risk of a collapse of competence: a workforce skilled at adopting AI tools but unable to maintain the systems they build.
@@
Early hypotheses and evidence from educational contexts have already started to question whether GenAI acts as a tool or is becoming a crutch (Balzan et al., 2026). According to a CDT national poll, 70% of teachers worry AI might weaken critical thinking and research skills, and more than half of students say they feel less connected to teachers when using AI tools (Laird et al., 2025). Surveys regarding ChatGPT-use, just a year after its unveiling, already raised fears that “learners may rely too heavily on the model” (Kasneci et al., 2023). Furthermore, “diminished epistemic vigilance, superficial learning and emotional dependence on AI” (Yan et al., 2025), as well as “metacognitive laziness” (Fan et al., 2025), have been observed in specific studies. Indeed, students themselves report cognitive offloading and over-reliance on GenAI tools (Wei et al., 2025). At the neural level, Kosmyna et al. (Kosmyna et al., 2025) provide electroencephalography (EEG) evidence: LLM-assisted essay writers exhibit the weakest brain network connectivity of all tested conditions and score lowest on ownership of their own writing, consistent with the cognitive outsourcing hypothesis. Lodge and Loble (Lodge and Loble, 2026) synthesize this evidence as a performance paradox: AI inflates short-term task performance while hollowing out the durable knowledge that is the true goal of learning. Beyond cognition, they argue that education is also a process of formation, the development of a thinking practitioner through the productive struggle of genuine inquiry. Unstructured AI, by removing this struggle, risks arresting that formation entirely.
A useful theoretical lens for this pattern is Kirschner’s (Kirschner, 2026) distinction between cognitive offloading and cognitive outsourcing, grounded in decades of cognitive and learning science research (Flavell, 1979). This framing motivates our central empirical question: does unrestricted GenAI use lead to the accumulation of epistemic debt, and if so, can structured intervention prevent it?
@@
Can we design AI interactions and interventions that allow novices to leverage the speed and productivity benefits of GenAI tools without sacrificing the depth of their learning? In other words, can high functional utility and genuine corrective competence coexist?
@@
To that end, we draw on the Teachable Agent paradigm (Biswas et al., 2005) to design an Explanation Gate intervention: before accepting AI-generated code, learners must explain its causal logic to the system in their own words. We hypothesize that this introduction of metacognitive friction can restore learner comprehension without significantly degrading productivity.
@@
  • RQ1 (Functional Utility & Velocity): To what extent does the introduction of automated metacognitive friction (the Explanation Gate) impact the immediate functional utility and development velocity of novice programmers compared to unrestricted GenAI usage?
@@
  • RQ3 (Interactional Stances): What qualitative differences in human-AI interaction strategies (e.g., consultant vs. contractor stances) correlate with the retention of corrective competence in unrestricted GenAI environments?
@@
  4. (4)
    Open Research Artifact: We release the complete VibeCheck extension, the Course Scheduler task suite (starter scaffold, reference implementation, and logic bomb version), the 12-assertion grading harness, and the full three-condition study replication protocol as an open artifact. [^2] https://github.com/sreecharansankaranarayanan/vibecheck
The following section grounds this argument in three complementary learning science frameworks.
@@ 2. Theoretical Framework
Our investigation is grounded in a synthesis of three complementary theories of learning: Cognitive Load Theory (CLT) (Sweller, 1988), the Script Theory of Guidance (SToG) (Fischer et al., 2013), and the Desirable Difficulties framework (Bjork, 1994). We argue that the interaction between novice programmers and Large Language Models (LLMs) represents a novel pedagogical context where traditional scaffolding metaphors break down, requiring a re-evaluation of how automated assistance supports, or subverts, schema acquisition.
To understand the consequences of unrestricted AI scaffolding, it is vital to distinguish epistemic debt from traditional technical debt. While technical lag often accumulates passively as systems age (Kruchten et al., 2012), standard technical debt is generally understood as a conscious trade-off: sacrificing long-term architecture for short-term delivery. Epistemic debt, conversely, is an invisible accumulation of unearned code. In novice programming, Generative AI accelerates this divergence. Because AI models lack “metacognitive calibration” and project unwarranted certainty (Quattrociocchi et al., 2025), novices absorb this forced confidence. They bypass the necessary cognitive friction required to construct accurate mental models, resulting in fragile experts who possess high functional utility but critically low corrective competence.
@@
One critical error in current EdTech discourse is the conflation of efficiency (productivity) with learning (schema construction). To dismantle this, we rely on CLT (Sweller, 1988), specifically the interplay between three types of load:
@@
    Germane Load: The mental effort dedicated to processing information into long-term memory schemas.
Expert developers use tools to minimize extraneous load, freeing up working memory for high-level problem solving (germane load). Kirschner (Kirschner, 2026) defines this as cognitive offloading. However, novices lack the schemas to distinguish between syntax (extraneous) and logic (intrinsic). When a novice uses GenAI to vibe code, the tool handles both the syntax and the underlying logic. This transition from offloading to cognitive outsourcing allows the learner to bypass the intrinsic load entirely (Kirschner, 2026). Without engaging with intrinsic load, germane processing cannot occur, and the learner accumulates epistemic debt. Recent empirical evidence from Yan et al. (Yan et al., 2026) suggests that this outsourcing does not merely bypass effort, but fundamentally disrupts the learner’s sense of agency and epistemic ownership. Without this ownership, the developer loses the “epistemic vigilance” required to catch the abstraction leaks that characterize vibe coding. This prediction finds experimental support in Shen and Tamkin (Shen and Tamkin, 2026), whose randomized study of professional developers learning a new asynchronous programming library found a 17% reduction in conceptual understanding, code reading, and debugging ability in the AI-assisted condition ($d=0.74$, $p=.01$), with no statistically significant gain in task completion speed.
@@ 2.2. Overscripting
SToG posits that external scripts (prompts, roles, instructions) are necessary to help learners coordinate complex tasks (Fischer et al., 2013). Ideally, these scripts provide a temporary structure that is gradually “faded” as the learner’s internal scripts (schemas) mature (Wecker and Fischer, 2010).
We argue that the “Copy-Paste” AI workflow (or worse, the one-click “Accept Changes”) represents a pathological case of overscripting (Dillenbourg, 2002). Because standard LLMs provide complete, high-fidelity solutions instantly, they act as a rigid script that replaces, rather than regulates, the learner’s cognitive activity. Unlike human tutors who withhold answers to prompt reflection, an unrestricted LLM resolves the problem space immediately, denying the learner the “productive failure” necessary for transfer (Kapur, 2008).
To counteract this, we introduce the concept of metacognitive friction (Flavell, 1979): a design paradigm that deliberately slows down the interaction loop to force generative processing. Operationally, we implement this through a metacognitive script in the SToG sense: a structured external prompt that guides the learner through the cognitive steps the AI would otherwise bypass. This aligns with Bjork’s concept of desirable difficulties (Bjork, 1994), specifically the generation effect, which posits that long-term retention is improved when learners must actively retrieve and reformulate knowledge rather than passively consuming it.
@@ 2.3. Scalable Assessment as the Gatekeeper
Historically, the Teachable Agent paradigm, where students learn by teaching a computer agent, has proven highly effective for deeper learning (Biswas et al., 2005). However, verifying that a student has genuinely “taught” the agent, rather than merely gaming the system (Baker, 2005), has traditionally required human Teaching Assistants (TAs), severely limiting the scalability of such interventions.
Recent advancements in LLM-based assessment allow us to overcome this bottleneck. By leveraging an “LLM-as-a-Judge” framework (Zheng et al., 2023; Choma et al., 2025), we adapt the Teachable Agent model into an automated, real-time Explanation Gate driven by metacognitive scripts. This mechanism deliberately reintroduces cognitive friction into the AI-assisted workflow to mitigate the rapid accumulation of epistemic debt. It requires the learner to articulate the underlying causal logic of the AI-generated code before integration is permitted, a direct operationalization of the self-explanation effect (Chi et al., 1989), which demonstrates that generating explanations during learning produces substantially deeper encoding than passive study. By acting as a semantic validity check, the Explanation Gate closes the loop on cognitive outsourcing by forcing novices to construct their own internal representations, ensuring that the student’s cognitive grasp scales concurrently with the codebase’s complexity.
With this theoretical scaffold in place, we describe the experimental design that operationalizes these principles.
@@ 3.2. The Coding Environment
  • Plugin Architecture: We developed a custom VS Code extension, VibeCheck  [^3] https://github.com/sreecharansankaranarayanan/vibecheck, which acted as a middleman between the LLM Sidebar and the editor. As illustrated in Figure 1, the extension implemented two core logic layers:
@@
      Rule Enforcement Engine: A background listener that detected large AI-generated insertions by monitoring document-change events. Insertions of two or more lines or 50 or more characters atomically were flagged as likely AI-origin and routed to the Explanation Gate Manager. Checkpoint state was enforced through a layered defense: save-interception, a post-save fallback, and a file-system watcher that caught disk-level bypass attempts (e.g., Cursor’s “Keep File” action).
@@
      Explanation Gate Manager (the friction loop): A state machine responsible for the intervention. Upon receiving a halt signal from the Enforcement Engine, this manager rendered a blocking modal overlay (Figure 3). It orchestrated an asynchronous loop with the Judge API, preventing the code merge until the Judge returned a “Pass” evaluation, closing the feedback loop shown in the diagram. Direct editor bypass attempts (editing the file without using the Apply button) were also intercepted and reverted automatically.
@@
      Telemetry Logger: Recorded gate-encounter metadata with millisecond-precision timestamps: gate triggers, explanation submissions (length only, never content), Judge scores, pass/fail outcomes, and total attempts per encounter. No code snippets or explanation text were stored (privacy-first design; all data remained local to the participant’s machine).